OpenStudy (anonymous):

Solve, using linear combination. 5x + 2y = -4 -3x + 2y = 12 (Points : 4) (-2, 3) (0, -2) (2, -7) (4, 0)

OpenStudy (anonymous):

2,-7

OpenStudy (anonymous):

how

OpenStudy (anonymous):

Isnt it 2,-3

OpenStudy (anonymous):

-2,3

OpenStudy (anonymous):

they gave u the answers already so all i did was i try all of them till i get the answers -4 and 12

OpenStudy (anonymous):

Ok thanks

OpenStudy (anonymous):

Can you guys help me on few more

OpenStudy (anonymous):

THis is the graph with the problem

OpenStudy (anonymous):

Is it -3,0

OpenStudy (anonymous):

2)0,3 cuz it didn't move its stay on zero and then to the right the point is 3

OpenStudy (anonymous):

First, multiply the first one by -1 5x + 2y = -4 -3x + 2y = 12 -5x-2y=4 -3x+2y=12 Notice the -2y and +2y. -5x-2y=4 -3x+2y=12 ------------ -8x =16 -8x=16 /-8 /-8 x= -2 Now, substitute x in the first equation to find y: 5(-2) + 2y = -4 -10+2y=-4 +10 +10 2y=6 /2 /2 y=3 (-2, 3)

OpenStudy (anonymous):

I already did that one got the same answer to i am talking about the graph

OpenStudy (anonymous):

is it -3,0

OpenStudy (anonymous):

talking about the graph that i attached

OpenStudy (anonymous):

Or 3,0

OpenStudy (anonymous):

Yes, the point is at (-3, 0) so I'd assume so. :o) And x-y=3 x+y=3 ------ x=3 3-y=3 -3 -3 y=0

OpenStudy (anonymous):

No, the points at positive 3, sorry

OpenStudy (anonymous):

its in the right side so its positive

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

So is it 3,0 or -3,0

OpenStudy (anonymous):

(3, 0) :o)

OpenStudy (anonymous):

3,0 is the answer right cause i got that answer to

OpenStudy (anonymous):

I think so :o)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok cool can you help me on a few more

OpenStudy (anonymous):

How many solutions are there to the following system of equations? 4x - 14y = 6 -2x + 7y = -3 (Points : 4) 0 1 2 infinitely many

OpenStudy (anonymous):

Dont worry i am trying to figure it out to so you wont feel i am taking advantage

OpenStudy (anonymous):

Is it 0 solutions

OpenStudy (anonymous):

I'm not sure how to do this one. Maybe this page can help you? :o) http://answers.yahoo.com/question/index?qid=20100108140909AAuJUwN

OpenStudy (anonymous):

I already got the answer i am going with 0 solutions

OpenStudy (anonymous):

its 0

OpenStudy (anonymous):

Solve this system of equations using any method. 3x - y = 4 x + 2y = 13 (Points : 4) (-1, 7) (4, 8) (5, 11) (3, 5)

OpenStudy (anonymous):

I caculated it is 3,5

OpenStudy (anonymous):

Do you want help with it then?

OpenStudy (anonymous):

Witch problem

OpenStudy (anonymous):

3x - y = 4 x + 2y = 13

OpenStudy (anonymous):

Is it 3,5 i just plug it in

OpenStudy (anonymous):

And it equaled it

OpenStudy (anonymous):

Ok.

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

cool few more please!

OpenStudy (anonymous):

Its not a graph i need help O:

OpenStudy (anonymous):

Is it Option B

OpenStudy (anonymous):

Are you guys there

OpenStudy (anonymous):

im here,im trying but i keep on getting B

OpenStudy (anonymous):

Me to

OpenStudy (anonymous):

then im guessing thats the answer,hold on i tell u in a minute my sister trying she have an a in math while i have a b

OpenStudy (anonymous):

she said B is the answer

OpenStudy (anonymous):

Lol tell her thanks so much

OpenStudy (anonymous):

I knew it

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

can you girls/Woman help me on a few more about 3

OpenStudy (anonymous):

Each child ticket for a ride costs $3, while each adult ticket costs $5. If the ride collected a total of $115, and 33 tickets were sold, how many of each type of ticket was sold? (Points : 4) 8 children and 25 adults 17 children and 16 adults 156 children and 41 adults 8 adults and 25 children

OpenStudy (anonymous):

sure im a junior,shes a senior

OpenStudy (anonymous):

I am 90% sure that 8 children and 25 adults

OpenStudy (anonymous):

Is the answer

OpenStudy (anonymous):

Or 8 adults and 25 children

OpenStudy (anonymous):

its 8adults and 25 children

OpenStudy (anonymous):

Dam flabbit

OpenStudy (anonymous):

How many solutions are there to the following system of equations? 2x - y = 2 -x + 5y = 3 (Points : 4) 0 1 2 infinitely many

OpenStudy (anonymous):

I knew it was either one of those XD

OpenStudy (anonymous):

I think there 1 solution

OpenStudy (anonymous):

1

OpenStudy (anonymous):

Yes i am correct!

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

THis one is a hand full number 8 is the one hopefully we both are going to work out

OpenStudy (anonymous):

Is it option A

OpenStudy (anonymous):

The top answer

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

a

OpenStudy (anonymous):

Sweet

OpenStudy (anonymous):

Solve the following system of equations. x + y = 3 2x + 2y = 6 (Points : 4) (0, 3) (1.5, 1.5) no solution infinite number of solutions

OpenStudy (anonymous):

Hmm

OpenStudy (anonymous):

I think its 2 solutions

OpenStudy (anonymous):

So i am going with infinite number of solutions

OpenStudy (anonymous):

that what i got 2

OpenStudy (anonymous):

put D

OpenStudy (anonymous):

You sure i feel shocked

OpenStudy (anonymous):

O:

OpenStudy (anonymous):

let me ask my sister

OpenStudy (anonymous):

Cause i usualy ok with these problems i am just shocked we got the same answer

OpenStudy (anonymous):

You guys there still

OpenStudy (anonymous):

yea,she said hold on

OpenStudy (anonymous):

k good with me

OpenStudy (anonymous):

its 2

OpenStudy (anonymous):

So infinite

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Which equation will result if the substitution method is used to solve the following system of linear equations? x + y = 6 x = y - 2 (Points : 4) (y - 2) + y = 6 x + y = y - 2 x + (y -2) = 6 x = (x- 6) - 2

OpenStudy (anonymous):

This is the new one i just need this one and one more then i will give you best response and a medal

OpenStudy (anonymous):

I think it is

OpenStudy (anonymous):

x + y = y - 2

OpenStudy (anonymous):

Witch is B

OpenStudy (anonymous):

i need my sister again,hold on

OpenStudy (anonymous):

Ok tell her to use this one cause i double check my asnwer and i got this